OpenStudy (anonymous):

What is the slope-intercept of this equation? 8x + 6y = 12 Please help me I really need the answer I'm doing a test

OpenStudy (kirbykirby):

\[ 8x + 6y = 12\\ 6y=-8x+12\\ y=-\frac{8}{6}x+\frac{12}{6}\\ y=-\frac{4}{3}x+2\]